OpenStudy (anonymous):

Find the coefficient of x^8 in (x^4 +x^5 + x^6)(x^3 +x^4 +x^5) (1+x+x^2+x^3+.....) Use generating function Please, help

OpenStudy (anonymous):

any idea, friend

OpenStudy (amistre64):

youd have to teach me what a generating function is

OpenStudy (anonymous):

lol... don't tease me, please

OpenStudy (amistre64):

im sure i can come up with a coeff, but with a different method

OpenStudy (anonymous):

ok, tell me yours

OpenStudy (amistre64):

x^4 +x^5 + x^6 x^3 +x^4 +x^5 ---------------- x^7+x^8+x^9 x^8+x^9+x^10 x^9+x^10+x^11 -------------------------- x^7 + 2x^8 + other stuff that will go above x^8

OpenStudy (amistre64):

1+x+x^2+x^3+..... x^7+2x^8 -------------- x^7+x^8+x^9+x^10+..... +2x^8+2x^9+... ------------------- 3x^8

OpenStudy (anonymous):

got it, friend

OpenStudy (anonymous):

thanks a lot. sorry, I didn't realize that you got the final answer. Lol .

OpenStudy (anonymous):

one more question. if jx^7/1-x =\[\sum_{k=0}^{\infty}\left(\begin{matrix}k+1 \\ 1\end{matrix}\right)x^k+7\]

OpenStudy (anonymous):

the last one is x^(k+7)

OpenStudy (anonymous):

sorry, my bad computer. x^7 / 1-x = sum (k+1 choose 1) of x^(k+7)